# Kronecker-product random matrices and a matrix least squares problem

### v1

@@ 1. Introduction
Here, $\boldsymbol{\theta},\boldsymbol{\xi},\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ are deterministic ridge-regularization and linear constraint parameters, and we will study a setting of random inputs given by independent Wigner matrices $A,B\in\mathbb{R}^{n\times n}$. Notably, the optimization variable $X$ has dimension comparable to $A$ and $B$. This problem (1) may be written equivalently in terms of the vectorization $\mathbf{x}=\operatorname{vec}(X)\in\mathbb{R}^{n^{2}}$ and diagonal matrices $\Theta=\operatorname{diag}(\boldsymbol{\theta})$ and $\Xi=\operatorname{diag}(\boldsymbol{\xi})$, as
@@
where $G_{0}\in\mathbb{C}^{n^{2}\times n^{2}}$ is a diagonal deterministic-equivalent matrix. Some high-level aspects of this model that may be suggestive of properties to be expected also in nonlinear models such as (2) and (3) include a characterization of the large-$n$ limit by a pair of fixed-point equations in an operator algebra rather than over the scalar field, and a “two-tiered” mean-field structure as reflected in (6), which arises from separate mean-field approximations over $A$ and $B$.
@@ 1.1. Proof ideas
$G_{ii}=\Big{(}a_{ii}I+B+\theta_{i}\Xi-zI-\sum_{r,s}^{(i)}a_{ir}G_{rs}^{(i)}a_{si}\Big{)}^{-1},\qquad G_{ij}={-}G_{ii}\sum_{r}^{(i)}a_{ir}G_{rj}^{(i)}.$ (9)
Applying concentration of the quadratic form in the first expression and averaging over $i=1,\ldots,n$, we will eventually obtain an approximate fixed-point relation for the partial trace
@@ 2.1. Kronecker deformed Wigner model
    $\mathcal{A}$ is a von Neumann algebra with unit $1_{\mathcal{A}}$, operator norm $\|\cdot\|_{\mathrm{op}}$, and faithful, normal, tracial state $\tau:\mathcal{A}\to\mathbb{C}$.
@@
We remark that $G_{0}(z)$ is a deterministic diagonal matrix in $\mathbb{C}^{n^{2}\times n^{2}}$. (We refer to [32, Theorem 4.4] for a previous example of this type of construction, in a different model.)
In contrast to more classical random matrix models, the above Stieltjes transform $m_{0}(z)$ in general does not seem to admit a simple characterization in terms of scalar-valued fixed-point equations. However, it may be characterized via a pair of fixed-point equations in the operator algebra $\mathcal{A}$, and $G_{0}(z)$ may be constructed from this characterization in the following way.
@@
In the setting of Assumption 2.2, set $\mathcal{A}^{+}=\{\mathsf{x}\in\mathcal{A}:\Im\mathsf{x}\geq\epsilon\text{ forsome }\epsilon>0\}$. For any $z\in\mathbb{C}^{+}$, there exist unique elements $\mathsf{m}_{a}(z),\mathsf{m}_{b}(z)\in\mathcal{A}^{+}$ satisfying the fixed-point equations
$\mathsf{m}_{a}(z)$ $=\frac{1}{n}\sum_{\alpha=1}^{n}(\mathsf{a}+\xi_{\alpha}\Theta-z-\mathsf{m}_{a}(z))^{-1},$ (19)
@@
Our main results for this model are the following estimates for its Stieltjes transform $m(z)$ and resolvent $G(z)$. Here, the stochastic domination notation $\xi\prec\zeta$ (c.f. [22]) means $\mathbb{P}[|\xi|>n^{\epsilon}\zeta]\leq n^{-D}$ for any fixed $\epsilon,D>0$ and all $n\geq n_{0}(\epsilon,D)$; we review this definition in Section 3.1.
@@
The estimates (25), (26), and (27) indicate that the entries of the resolvent fall on three scales of orders 1, $n^{-1/2}$, and $n^{-1}$, as depicted in the numerical simulation of Figure 1.
@@ 3. Preliminaries
In the remainder of the paper, we prove the preceding results. We summarize in this section the technical tools needed for our analyses. Section 4 contains the core of our main argument for analyzing a deformed Wigner-type matrix model in an operator-algebra setting. Section 5 completes the analyses for the Kronecker deformed Wigner model (17), and Section 6 completes the analyses for the least-squares problem (30).
@@
and $\phi(\sup\mathsf{x}_{i})=\lim\phi(\mathsf{x}_{i})$ for any bounded increasing net $\{\mathsf{x}_{i}\}$ of elements $\mathsf{x}_{i}\geq 0$ in $\mathcal{X}$. We will be working with $\mathcal{X}$-valued random variables, understood as strongly (i.e. Bochner) measurable functions from the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$ to the Banach space $(\mathcal{X},\|\cdot\|_{\mathrm{op}})$. For a $\mathcal{X}$-valued random variable $\mathsf{x}$ satisfying $\mathbb{E}\|\mathsf{x}\|_{\mathrm{op}}<\infty$, we denote by $\mathbb{E}\mathsf{x}\in\mathcal{X}$ its expectation and $\mathbb{E}[\mathsf{x}\mid\mathscr{G}]$ its conditional expectation with respect to a sub-sigma-field $\mathscr{G}\subset\mathscr{F}$ (c.f. [57, Chapter 1]).
@@ 3.1. Stochastic domination
The argument is similar to the scalar setting (c.f. [28, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.3),
@@
The triangle inequality for $\|\cdot\|_{p}$ implies for $\lambda\in[0,1]$ that $\|\lambda\mathsf{x}+(1-\lambda)\mathsf{y}\|_{p}\leq\lambda\|\mathsf{x}\|_{p}+(1-\lambda)\|\mathsf{y}\|_{p}$, so $\mathsf{x}\mapsto\|\mathsf{x}\|_{p}$ is continuous and convex. Then (c.f. [57, Proposition 1.12])
@@ 3.2. Khintchine-type inequalities
The following statements may be deduced from the non-commutative Rosenthal inequalities of [38]. In the setting of Rademacher variables, similar Khintchine inequalities have been shown in [56, Eq. (8.4.11)] and [59, Theorem 6.22].
@@ 3.3. Fluctuation averaging
The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [23, Theorem 4.7] and [29, Lemma A.2].
@@ 3.4. Minors and resolvent identities
$R=(H\otimes\mathsf{x}-\mathsf{z})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X},$
which exists by Lemma D.1 as ${-}\Im(H\otimes\mathsf{x}-\mathsf{z})=\Im\mathsf{z}\geq\epsilon$ if $\Im\mathsf{z}_{i}\geq\epsilon$ for every $i$. For an index set $S\subseteq\{1,\ldots,n\}$, we define $H^{(S)}\in\mathbb{C}^{n\times n}$ and $R^{(S)}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ by
@@
which are purely algebraic and identical to those of the scalar setting, see e.g. [26, Lemma 4.2]. ∎
@@ 3.5. Maximum modulus principle
We will use the following quantitative version of the maximum modulus principle, following [60, Appendix A].
@@
is increasing and convex [35]. Fixing $a>0$, consider the conformal mapping $\psi_{a}:\mathbb{D}\rightarrow\mathbb{C}^{+}$ given by $\psi_{a}(z)=ia\frac{1+z}{1-z}$. For any $z\in\mathbb{D}$,
@@ 4.1. Operator fixed-point equations
For existence and uniqueness in part (a), we apply the general result of [37, Theorem 2.1]: Let $\mathcal{B}$ be any $C^{*}$-algebra and denote its right operator half-plane
@@
To show that this solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$, we claim that the map $\mathsf{s}\mapsto I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$ is the $\mathcal{D}\otimes\mathcal{X}$-valued R-transform of $\mathsf{h}\otimes\mathsf{x}$, and (48) is the Cauchy-R relation over $\mathcal{D}\otimes\mathcal{X}$. We recall a few relevant notions from operator-valued free probability theory [47, Chapter 9.2]: Let $\operatorname{NC}(m)$ be the lattice of non-crossing partitions of $\{1,\ldots,m\}$. Associated to any von Neumann subalgebra $\mathcal{B}$ of $\mathcal{Y}=\mathcal{A}\otimes\mathcal{X}$ and its conditional expectation $\tau^{\mathcal{B}}:\mathcal{Y}\to\mathcal{B}$ (c.f. Lemma D.2) is a system of $\mathcal{B}$-valued cumulants $(\kappa_{\pi}^{\mathcal{B}})_{\pi\in\operatorname{NC}(m)}$, which are $\mathbb{C}$-multilinear maps $\kappa_{\pi}^{\mathcal{B}}:\mathcal{Y}^{m}\to\mathcal{B}$ satisfying the free moment-cumulant relation
and the bimodule properties
Specializing to our setting with $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$, we claim that for any $\mathsf{s}\in\mathcal{B}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
@@
Indeed, since $\mathsf{h}$ is free of $\mathcal{M}$ and hence of $\mathcal{D}\subset\mathcal{M}$, it is readily checked by definition (c.f. [47, Definition 9.5]) that $\mathsf{h}\otimes\mathsf{x}$ is free of $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$ with amalgamation over $1\otimes\mathcal{X}$ under the conditional expectation $\tau\otimes 1:\mathcal{A}\otimes\mathcal{X}\to\mathcal{X}\cong 1\otimes\mathcal{X}$. Then by [50, Theorem 3.6], for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
@@
where $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$ are the scalar-valued free cumulants of $\mathsf{h}$: This may be verified by expressing $\kappa_{m}^{1\otimes\mathcal{X}}(\mathsf{h}\otimes 1,\ldots,\mathsf{h}\otimes 1)$ in terms of moments under the conditional expectation $\tau\otimes 1$ using Möbius-inversion of (50) (c.f. [47, Eq. (9.19)]), applying the identity $(\tau\otimes 1)[\mathsf{a}\otimes 1]=\tau(\mathsf{a})(1\otimes 1)$ for each moment term, and re-applying (50) to express the result back in terms of the scalar cumulants $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$. Here, $\mathsf{h}\in\mathcal{A}$ is semi-circular, so $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})=1$ if $m=2$ and 0 otherwise (c.f. [51, Example 11.21]). Thus only the $m=2$ term above remains, and we obtain the claim (53).
@@
where we used $L^{p}$-contractivity of $\tau^{\mathcal{D}}\otimes 1$ (Lemma D.4), Hölder’s inequality (Lemma D.3), the given conditions $\Im\mathsf{z}\geq\delta$ and $\Im(\mathsf{z}+\Delta)\geq\delta/2$, and Lemma D.1.
@@
Here, the second line uses Hölder’s inequality (Lemma D.3), $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$, contractivity of the conditional expectation $\tau\otimes 1$ in $L^{p}$ (Lemma D.4), and $\|\mathsf{r}\|_{\mathrm{op}}\leq\delta^{-1}$ by the assumption $\Im\mathsf{z}\geq\delta$ and Lemma D.1. On the other hand, differentiating (56) at $\omega_{0}=\omega(0)$ and using $({-}\mathsf{z}_{i}-\mathsf{x}\omega_{0}\mathsf{x})^{-1}=({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}=(\mathsf{r}_{0})_{ii}$ by Lemma 4.3(a), we have

### v2

@@ 1. Introduction
Here, $\boldsymbol{\theta},\boldsymbol{\xi},\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ are deterministic ridge-regularization and linear constraint parameters, and we will study a setting of random inputs given by independent Wigner matrices $A,B\in\mathbb{R}^{n\times n}$. Notably, the optimization variable $X=(x_{ij})_{i,j=1}^{n}$ (with $x_{ij}$ denoting the entries of $X$) has dimension comparable to $A$ and $B$. This problem (1) may be written equivalently in terms of the vectorization $\mathbf{x}=\operatorname{vec}(X)\in\mathbb{R}^{n^{2}}$ and diagonal matrices $\Theta=\operatorname{diag}(\boldsymbol{\theta})$ and $\Xi=\operatorname{diag}(\boldsymbol{\xi})$, as
@@
where $G_{0}\in\mathbb{C}^{n^{2}\times n^{2}}$ is a diagonal deterministic-equivalent matrix. We remark that although this model (1) is a “toy” setting that simplifies many of the additional complexities of nonlinear models such as (2) and (3), some high-level aspects of this model may be suggestive of properties to be expected also in these nonlinear models, including a characterization of the large-$n$ limit by a pair of fixed-point equations in an operator algebra rather than over the scalar field, and a “two-tiered” mean-field structure as reflected in (6), which arises from separate mean-field approximations over $A$ and $B$.
@@ 1.1. Proof ideas
$G_{ii}=\Big{(}a_{ii}I+B+\theta_{i}\Xi-zI-\sum_{r,s}^{(i)}a_{ir}G_{rs}^{(i)}a_{si}\Big{)}^{-1},\qquad G_{ij}={-}G_{ii}\sum_{r}^{(i)}a_{ir}G_{rj}^{(i)},$ (9)
where $G_{rs}^{(i)}$ is the version of $G_{rs}$ setting to 0 the $i^{\text{th}}$ row and column of $A$, and $\sum_{r}^{(i)}$ is the summation over indices $r\neq i$. Applying concentration of the quadratic form in the first expression and averaging over $i=1,\ldots,n$, we will eventually obtain an approximate fixed-point relation for the partial trace
@@ 2.1. Kronecker deformed Wigner model
    $\mathcal{A}$ is a von Neumann algebra with unit $1_{\mathcal{A}}$, operator norm $\|\cdot\|_{\mathrm{op}}$, and faithful, normal, tracial state $\tau:\mathcal{A}\to\mathbb{C}$. (We review this definition at the start of Section 3.)
@@
We remark that $G_{0}(z)$ is a deterministic diagonal matrix in $\mathbb{C}^{n^{2}\times n^{2}}$. (We refer to [33, Theorem 4.4] for a previous example of this type of construction, in a different model.)
In contrast to more classical random matrix models, the above Stieltjes transform $m_{0}(z)$ in general does not seem to admit a simple characterization in terms of scalar-valued fixed-point equations. In some previously studied models with variance profiles or correlated entries, deterministic equivalents for the Stietljes transform and resolvent may instead be defined via fixed-point equations in a vector or matrix space [5, 26, 4]. For our model of interest, the most direct characterization of its asymptotic Stieltjes transform seems to be via the following fixed-point equation in the operator algebra $\mathcal{A}$, from which a deterministic equivalent $G_{0}(z)$ for its resolvent may also be constructed.
@@
In the setting of Assumption 2.2, set $\mathcal{A}^{+}=\{\mathsf{x}\in\mathcal{A}:\Im\mathsf{x}\geq\epsilon\text{ forsome }\epsilon>0\}$. For any $z\in\mathbb{C}^{+}$, there exists a unique element $\mathsf{m}_{a}(z)\in\mathcal{A}^{+}$ satisfying the fixed-point equation
$\mathsf{m}_{a}(z)$ $=\frac{1}{n}\sum_{\alpha=1}^{n}(\mathsf{a}+\xi_{\alpha}\Theta-z-\mathsf{m}_{a}(z))^{-1}.$ (19)
Similarly, there exists a unique element $\mathsf{m}_{b}(z)\in\mathcal{A}^{+}$ satisfying
@@
Qualitative approximations of the resolvent $G(z)$ and Stieltjes transform $m(z)$ by the above constructions $G_{0}(z)$ and $m_{0}(z)$ may be obtained using standard techniques of free probability theory, such as a moment expansion of $G(z)$ and $m(z)$ for large $\Im z$ and analytic continuation to the full upper half-plane. A primary purpose of our work is to obtain more quantitative estimates that are not readily deduced from these types of methods. The following theorem states these estimates, which represent optimal approximations for $G(z)$ and $m(z)$ on the global spectral scale. Here, the stochastic domination notation $\xi\prec\zeta$ (c.f. [23]) means $\mathbb{P}[|\xi|>n^{\epsilon}\zeta]\leq n^{-D}$ for any fixed $\epsilon,D>0$ and all $n\geq n_{0}(\epsilon,D)$; we review this definition in Section 3.1.
@@
The estimates (25), (26), and (27) indicate that the entries of the resolvent fall on three scales of orders 1, $n^{-1/2}$, and $n^{-1}$, as depicted in the numerical simulation of Figure 1. In a setting where $\Theta\otimes\Xi$ commutes with $A\otimes I$ and $I\otimes B$, the origins and optimality of these estimates may be understood via a contour integral representation of $G(z)$, as we discuss in Appendix A. The above theorem illustrates that various properties of $G(z)$ for general $\Theta\otimes\Xi$ are in fact similar to this commutative case.
@@ 3. Preliminaries
In the remainder of the paper, we prove the preceding results. We note that the bulk of our analysis is in fact carried out for a generalized resolvent operator
$R=\bigg{(}H\otimes\mathsf{x}-\sum_{k=1}^{K}D_{k}\otimes\mathsf{x}_{k}\bigg{)}^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$
in the setting of an abstract von Neumann algebra $\mathcal{X}$, which we define more precisely in Section 4. This analysis may be of independent interest to models beyond the ones we consider in our work.
We summarize in this section the technical tools needed for our analyses. Section 4 contains the core of our main argument for analyzing the above generalized resolvent in an operator-algebra setting. Section 5 completes the analyses for the Kronecker deformed Wigner model (17), and Section 6 completes the analyses for the least-squares problem (30).
@@
and $\phi(\sup\mathsf{x}_{i})=\lim\phi(\mathsf{x}_{i})$ for any bounded increasing net $\{\mathsf{x}_{i}\}$ of elements $\mathsf{x}_{i}\geq 0$ in $\mathcal{X}$. We will be working with $\mathcal{X}$-valued random variables, understood as strongly (i.e. Bochner) measurable functions from the underlying probability space $(\Omega,\mathscr{F},\mathbb{P})$ to the Banach space $(\mathcal{X},\|\cdot\|_{\mathrm{op}})$. For a $\mathcal{X}$-valued random variable $\mathsf{x}$ satisfying $\mathbb{E}\|\mathsf{x}\|_{\mathrm{op}}<\infty$, we denote by $\mathbb{E}\mathsf{x}\in\mathcal{X}$ its expectation and $\mathbb{E}[\mathsf{x}\mid\mathscr{G}]$ its conditional expectation with respect to a sub-sigma-field $\mathscr{G}\subset\mathscr{F}$ (c.f. [58, Chapter 1]).
@@ 3.1. Stochastic domination
The argument is similar to the scalar setting (c.f. [29, Lemma D.2]): For each fixed $p\in[1,\infty)$, by the triangle inequality and Hölder’s inequality for the $L^{p}$-norm (Lemma D.2),
@@
The triangle inequality for $\|\cdot\|_{p}$ implies for $\lambda\in[0,1]$ that $\|\lambda\mathsf{x}+(1-\lambda)\mathsf{y}\|_{p}\leq\lambda\|\mathsf{x}\|_{p}+(1-\lambda)\|\mathsf{y}\|_{p}$, so $\mathsf{x}\mapsto\|\mathsf{x}\|_{p}$ is continuous and convex. Then (c.f. [58, Proposition 1.12])
@@ 3.2. Khintchine-type inequalities
The following statements may be deduced from the non-commutative Rosenthal inequalities of [39]. In the setting of Rademacher variables, similar Khintchine inequalities have been shown in [57, Eq. (8.4.11)] and [60, Theorem 6.22].
@@ 3.3. Fluctuation averaging
The following fluctuation averaging statements in the $L^{p}$-norms on $\mathcal{X}$ are similar to those of the scalar setting, see e.g. [24, Theorem 4.7] and [30, Lemma A.2].
@@ 3.4. Minors and resolvent identities
$R=(H\otimes\mathsf{x}-\mathsf{z})^{-1}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}.$
We note that this generalized resolvent exists by the following standard lemma, since $\Im(H\otimes\mathsf{x}-\mathsf{z})={-}\Im\mathsf{z}\leq{-}\epsilon$.
Lemma 3.6 ([35], Lemma 3.1).
Suppose $\mathsf{x}\in\mathcal{X}$ is such that for some $\epsilon>0$, either $\Im\mathsf{x}\geq\epsilon$ or $\Im\mathsf{x}\leq-\epsilon$. Then $\mathsf{x}$ is invertible and $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq 1/\epsilon$.
For an index set $S\subseteq\{1,\ldots,n\}$, we define $H^{(S)}\in\mathbb{C}^{n\times n}$ and $R^{(S)}\in\mathbb{C}^{n\times n}\otimes\mathcal{X}$ by
@@
which are purely algebraic and identical to those of the scalar setting, see e.g. [27, Lemma 4.2]. ∎
@@ 3.5. Maximum modulus principle
We will use the following quantitative version of the maximum modulus principle, following [61, Appendix A].
@@
is increasing and convex [36, Section 1]. Fixing $a>0$, consider the conformal mapping $\psi_{a}:\mathbb{D}\rightarrow\mathbb{C}^{+}$ given by $\psi_{a}(z)=ia\frac{1+z}{1-z}$. For any $z\in\mathbb{D}$,
@@ 4.1. Operator fixed-point equations
In the following we recall a few relevant notions from operator-valued free probability theory [48, Chapter 9.2]: Let $\operatorname{NC}(m)$ be the lattice of non-crossing partitions of $\{1,\ldots,m\}$. Associated to any von Neumann subalgebra $\mathcal{B}$ of $\mathcal{Y}=\mathcal{A}\otimes\mathcal{X}$ and its conditional expectation $\tau^{\mathcal{B}}:\mathcal{Y}\to\mathcal{B}$ (c.f. Lemma D.1) is a system of $\mathcal{B}$-valued cumulants $(\kappa_{\pi}^{\mathcal{B}})_{\pi\in\operatorname{NC}(m)}$, which are $\mathbb{C}$-multilinear maps $\kappa_{\pi}^{\mathcal{B}}:\mathcal{Y}^{m}\to\mathcal{B}$ satisfying the free moment-cumulant relation
and the bimodule properties
@@
For existence and uniqueness in part (a), we apply the general result of [38, Theorem 2.1]: Let $\mathcal{B}$ be any $C^{*}$-algebra and denote its right operator half-plane
@@
To show that this solution is given by $\mathsf{r}_{0}=(\tau^{\mathcal{D}}\otimes 1)[(\mathsf{h}\otimes\mathsf{x}-\mathsf{z})^{-1}]$, we claim that the map $\mathsf{s}\mapsto I\otimes\mathsf{x}(n^{-1}\operatorname{Tr}\otimes 1[\mathsf{s}])\mathsf{x}$ is the $\mathcal{D}\otimes\mathcal{X}$-valued R-transform of $\mathsf{h}\otimes\mathsf{x}$, and (51) is the Cauchy-R relation over $\mathcal{D}\otimes\mathcal{X}$. Namely, for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
@@
Indeed, since $\mathsf{h}$ is free of $\mathcal{M}$ and hence of $\mathcal{D}\subset\mathcal{M}$, it is readily checked by definition (c.f. [48, Definition 9.5]) that $\mathsf{h}\otimes\mathsf{x}$ is free of $\mathcal{B}=\mathcal{D}\otimes\mathcal{X}$ with amalgamation over $1\otimes\mathcal{X}$ under the conditional expectation $\tau\otimes 1:\mathcal{A}\otimes\mathcal{X}\to\mathcal{X}\cong 1\otimes\mathcal{X}$. Then by [51, Theorem 3.6], for any $\mathsf{s}\in\mathcal{D}\otimes\mathcal{X}$ with $\|\mathsf{s}\|_{\mathrm{op}}$ small enough,
@@
where $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$ are the scalar-valued free cumulants of $\mathsf{h}$: This may be verified by expressing $\kappa_{m}^{1\otimes\mathcal{X}}(\mathsf{h}\otimes 1,\ldots,\mathsf{h}\otimes 1)$ in terms of moments under the conditional expectation $\tau\otimes 1$ using Möbius-inversion of (48) (c.f. [48, Eq. (9.19)]), applying the identity $(\tau\otimes 1)[\mathsf{a}\otimes 1]=\tau(\mathsf{a})(1\otimes 1)$ for each moment term, and re-applying (48) to express the result back in terms of the scalar cumulants $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})$. Here, $\mathsf{h}\in\mathcal{A}$ is semi-circular, so $\kappa_{m}(\mathsf{h},\ldots,\mathsf{h})=1$ if $m=2$ and 0 otherwise (c.f. [52, Example 11.21]). Thus only the $m=2$ term above remains, and we obtain the claim (53).
@@
where we used $L^{p}$-contractivity of $\tau^{\mathcal{D}}\otimes 1$ (Lemma D.3), Hölder’s inequality (Lemma D.2), the given conditions $\Im\mathsf{z}\geq\delta$ and $\Im(\mathsf{z}+\Delta)\geq\delta/2$, and Lemma 3.6.
@@
Here, the second line uses Hölder’s inequality (Lemma D.2), $\|\mathsf{x}^{-1}\|_{\mathrm{op}}\leq\gamma$, contractivity of the conditional expectation $\tau\otimes 1$ in $L^{p}$ (Lemma D.3), and $\|\mathsf{r}\|_{\mathrm{op}}\leq\delta^{-1}$ by the assumption $\Im\mathsf{z}\geq\delta$ and Lemma 3.6. On the other hand, differentiating (56) at $\omega_{0}=\omega(0)$ and using $({-}\mathsf{z}_{i}-\mathsf{x}\omega_{0}\mathsf{x})^{-1}=({-}\mathsf{z}_{i}-\mathsf{x}\mathsf{m}_{0}\mathsf{x})^{-1}=(\mathsf{r}_{0})_{ii}$ by Lemma 4.3(a), we have